THE RESISTIANCE OF WIRE IS R . ITS CUT INTO 4 EQUAL PARTS AND THESE PARTS ARE CONNECTED SIDE BY SIDE M THE EFFECTIVE RESISTANCE BECOMES?
1
Expert's answer
2019-09-27T09:57:59-0400

Resistance of single wire"=R"

When the wire is cut into four pieces,the resistance gets divided into one-fourth of the total resistance for each part.

So,resistance of each single part is "\frac{R}{4}"

When they are connected in series(side by side),their resistance gets add up;

Net resistance after connection in series"=\frac{R}{4}+\frac{R}{4}+\frac{R}{4}+\frac{R}{4}=\frac{4R}{4}=R"